OpenStudy (stephanos100):

Find an explicit rule for the nth term of the sequence. 7, -7, 7, -7, ... MEDALS!!!! GET YOUR MEDALS!!

OpenStudy (stephanos100):

is it an = 7 • (-1)n

OpenStudy (stephanos100):

as in times

OpenStudy (stephanos100):

a sub n = 7 * (-1)^n

jimthompson5910 (jim_thompson5910):

Did you mean this? \[\Large a_n = 7(-1)^{n}\]

OpenStudy (stephanos100):

yes

jimthompson5910 (jim_thompson5910):

oh nvm, you wrote `a sub n = 7 * (-1)^n` which is close. Normally n starts at n = 1 to indicate the first term n = 0 could be the starting point but then the count gets thrown off (eg: n = 3 refers to the fourth term, not the third, a bit confusing I know)

jimthompson5910 (jim_thompson5910):

since n = 1 is a more natural starting point, you have to change that 'n' to 'n-1' so the nth term formula is \[\Large a_n = 7(-1)^{n-1}\] where n starts at n = 1 and n is an integer

OpenStudy (stephanos100):

okay so thats the answer then?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (stephanos100):

and thats because when you substitute 1 for n it refers to the second number in the sequence right

jimthompson5910 (jim_thompson5910):

no, again it's more natural to say "n = 1 refers to the first term" Let's see what happens when we plug in n = 1 \[\Large a_n = 7(-1)^{n-1}\] \[\Large a_1 = 7(-1)^{1-1}\] \[\Large a_1 = 7(-1)^{0}\] \[\Large a_1 = 7(1)\] \[\Large a_1 = 7\] so 7 is our first term, which matches up with what the sequence gives us

jimthompson5910 (jim_thompson5910):

Let's try n = 2 \[\Large a_n = 7(-1)^{n-1}\] \[\Large a_2 = 7(-1)^{2-1}\] \[\Large a_2 = 7(-1)^{1}\] \[\Large a_2 = 7(-1)\] \[\Large a_2 = -7\] and we get -7 as our second term, as expected

OpenStudy (stephanos100):

oh okay thank you i got confused because i put that (-1)^0 equaled -1 and not 1

jimthompson5910 (jim_thompson5910):

I used the rule \[\Large x^0 = 1\] where x is not equal to zero and x can be equal to any other number

OpenStudy (stephanos100):

thank you!!!

jimthompson5910 (jim_thompson5910):

you're welcome